OpenStudy (anonymous):

Suppose the rate of growth of population is given by: dP/dT=0.010P(100-P), where time is in months and P(0) = 2 A) Find formula for the population in terms of T B) How long will it take the population to reach 60 individuals??

OpenStudy (loser66):

isolate dp take integral both sides to get P = ...... + C replace t = 0, P =2 to get C plug back to get the whole answer

OpenStudy (cruffo):

Not sure if you are familiar with the logistic differential equation: http://en.wikipedia.org/wiki/Logistic_function In particular, \[\frac{dP}{dt}=rP\left(1 - \frac{P}{K}\right)\] where the constant r defines the growth rate and K is the carrying capacity. This has the solution \[P(t) = \frac{K P_0 e^{rt}}{K + P_0 \left( e^{rt} - 1\right)} \] with \(P_0\) being the initial population.

OpenStudy (anonymous):

cam you shoew me step by step how to get the answer

OpenStudy (anonymous):

@cruffo @Loser66

OpenStudy (cruffo):

Using the formula above, or actually separating variables and integrating?

OpenStudy (anonymous):

using the formula

OpenStudy (cruffo):

Oh good, the integration would require use to find a partial fraction decomposition :) To use the formula we have to rewrite the right hand side to find out what r and K will be: Concentrate on the (100 - P) part first. We want this to look like (1 - P/K). Factor out 100 from the (100 - P) to get \(\;100(1 - \frac{P}{100})\). You can check the factoring by redistributing the 100. ok so far?

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Next step?

OpenStudy (cruffo):

Great, becaue that was the tricky part. To continue, we have \[0.01P(100 - P) = 0.01\cdot P \cdot 100 \cdot \left( 1 - \frac{p}{100}\right)\]

OpenStudy (anonymous):

Ok next?

OpenStudy (cruffo):

\[\frac{dP}{dt}=rP\left(1 - \frac{P}{K}\right)\] If you simplify the expression in the previous step, what do you have for r and K?

OpenStudy (anonymous):

multiple to cancel out?

OpenStudy (anonymous):

multiply

OpenStudy (cruffo):

right, \[ 0.01P(100 - P) = 0.01\cdot P \cdot 100 \cdot \left( 1 - \frac{p}{100}\right) = 1 P\left( 1 - \frac{p}{100}\right)\] so r = 1 and k = 100.

OpenStudy (anonymous):

next step???

OpenStudy (cruffo):

\[P(t) = \frac{K P_0 e^{rt}}{K + P_0 \left( e^{rt} - 1\right)}\] Replace K with 100, r with 1, and the initial population \(P_0 = P(0) = 2\).

OpenStudy (anonymous):

what about part B

OpenStudy (anonymous):

How long will it take for the population to reach 60 individuals?

OpenStudy (cruffo):

Here we have to solve the equation: \[\frac{200e^t}{100 + 2(e^t-1)} = 60\]

OpenStudy (anonymous):

multiply 60 with the den?

OpenStudy (cruffo):

Yep. That's what I would do :)

OpenStudy (anonymous):

6000+120e^t-120=200e^t

OpenStudy (anonymous):

5880=80e^t

OpenStudy (cruffo):

right!

OpenStudy (anonymous):

e^t=73.5 now what?

OpenStudy (cruffo):

take the natural log of both sides :)

OpenStudy (anonymous):

show me please!

OpenStudy (anonymous):

i wanna make sure i did it correctly

OpenStudy (cruffo):

\[e^t = 73.5\] \[\ln(e^t) = \ln(73.5)\] then use power rule : \[t \cdot \ln(e) = \ln(73.5)\] and ln(e) is just 1 so \[t = \ln(73.5)\]

OpenStudy (anonymous):

you my friend earned yourself a best response point!

OpenStudy (anonymous):

can you help me with one more question?

OpenStudy (cruffo):

yah, but make it a new post.